OpenStudy (anonymous):

find the domain and range h(x) =( 4-x^2)^.5

OpenStudy (anonymous):

since you can't have a negative square root, \[x ^{2} \le 4\] \[x \le 2\] and \[y \ge 0\]

OpenStudy (anonymous):

The domain is \(4-x^2\ge0 \). By testing its signs, we find that the domain is \([-2,2]\).

OpenStudy (anonymous):

range

OpenStudy (anonymous):

oops my mistake, thanks @Anwara for correcting !